OpenStudy (anonymous):

what is the distance in radical form between the points 4,6 8,-2

OpenStudy (mathstudent55):

You mean points (4, 6) and (8, -2)? If so, you need the distance formula between two points, (a, b) and (c, d) = sqrt( (a - c)^2 + (b - d)^2 ) In your case, sqrt( (4 - 8)^2 + ( 6 - (-2) )^2 ) = sqrt( (-4)^2 + 8^2 ) = sqrt(16 + 64) = sqrt(80) = sqrt(16*5) = 4sqrt(5)

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm